OpenStudy (anonymous):

Expand the following using the Binomial Theorem and Pascal’s triangle. (x + 2)^6

OpenStudy (isaiah.feynman):

Can you draw the pascal's triangle?

OpenStudy (anonymous):

I have it, but I don't understand it

OpenStudy (anonymous):

if the exponent is 6, does that mean I look on row 6?

OpenStudy (anonymous):

Yes Im pretty sure.

OpenStudy (anonymous):

Ok, I don't know what to do from there.

OpenStudy (anonymous):

Ok let me do the math 1 sec

OpenStudy (anonymous):

Hmm Im doing something wrong here i missed a step

OpenStudy (anonymous):

there's an example on my lesson:

OpenStudy (anonymous):

you have to expand it through row 2 and row 4

OpenStudy (anonymous):

what do u mean?

OpenStudy (anonymous):

let me see how to explain it.

OpenStudy (anonymous):

I get the part about the coefficients, but not the exponents. How do you know what numbers to put for the exponents

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

like if the give me 5 with the little exponent its like saying 5*5*5*5*5 *= times

OpenStudy (anonymous):

that's not what I asked.. never mind

OpenStudy (anonymous):

I found this:

OpenStudy (anonymous):

(a + b)^2 = (a + b)(a + b) = a^2 + ab + ba + b^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

so that's how we form it

OpenStudy (anonymous):

but there are 7 numbers in the last row.. not 3... so idk

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

multiply what?..

OpenStudy (anonymous):

@Isaiah.Feynman can you please help